Game 1 X Y 6,2 2,1 1,3 5,2 ,1 3,3 Z 3,3 C 4 1,22. In Game 1, if 0-2 = (1/4, 74, 12), what is BR2(0-2)?
Answer #1

2) Given O., - (44, 4/4, 12) we know Ons is probability of flayer 1s reshonse sil. A, B, C . So, the Game 1 :- 9 Player qailyd Mayed by player 2 expected outcome : - 15 (1*4) + (2x4) + (334) = 4+2+ = = 2.25 of Z is played by player 2 expected out& q, and %= o. e shf, Bh_ Ll.) = 9,91,1-9-90%. to, len, (0.) = (0,0,1))